OpenStudy (anonymous):

Write the equation of the line passing through the points (0, -4) and (-2, 2).

OpenStudy (anonymous):

First you have to find the slope which is (y2-y1), (x2-x1) 2-(-4)=6 -2-0=-2 so the slope is 6/-2 or just -3 Now that you have that, you have to substitute the x and the y variables for the x and y in the slope intercept equation say you use (0,-4) -4=-3(0)+b You would get b=-4 so the equation would be y=-3x-4 :D